# A Concentration Bound for TD(0) with Function Approximation

### v1

@@ 1 Introduction
TD(0) is one of the most popular reinforcement learning algorithms for policy evaluation (Tsitsiklis and Van Roy 1997). Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework. To mitigate the issues of large state spaces, the value function is often approximated using a linear combination of feature vectors. This algorithm is referred to as TD(0) with linear function approximation. In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain. Our goal in this paper is to obtain a bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
@@
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2023), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$). Korda and LA (2015) give a high probability bound for the same online framework as us.
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2021), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
Inspired by Chen et al. (2020), one of us considered stochastic approximation involving contractive maps and martingale noise, and derived such concentration bounds ‘from some $n_{0}$ on’ for this class of algorithms (Borkar 2021). This covered in particular synchronous Q-learning for discounted cost and some related schemes. In Chandak et al. (2022), we extended this work to cover ‘Markov noise’ (Meerkov 1972) in the stochastic approximation scheme, allowing us to give bounds for the asynchronous case. As mentioned before, this work assumed almost-sure boundedness of iterates, which is not satisfied by the TD(0) algorithm. We remove the need for this assumption in our current work. Other articles aiming at bounds of these forms, such as the one we provide, are found in Chandak et al. (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes.
@@ 2 TD(0) - Background and Main Result
The state space can often be large ($s\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}:S\mapsto\mathcal{R},1\leq i\leq d$, with $s>>d\geq 1$. Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$. Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\phi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$. Here $x$ denotes the learnable weights for the linear function approximator. Since $\{\phi_{i}\}$ are linearly independent, $\Phi$ is full rank. Substituting this approximation into the dynamic programming equation above leads to
@@
    First we make an assumption on $\Phi$. For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$. Assume that
@@
    Since the feature vectors can be scaled without affecting the algorithm (the weights $x(i)$ get scaled accordingly), this assumption does not restrict the algorithm. Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption.
@@
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. The proof for this lemma has been postponed to Appendix B.
@@
$\alpha=\sqrt{1-\min_{r\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
@@
The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$. Then,
@@
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
@@
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),0<\delta\leq C,$
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon),\delta>C.$
Here $C=\sqrt{d}c_{3}\left(2+\|x^{*}\|+\frac{a(n_{0})(c_{2}+1)+1}{1-\alpha-a(n_{0})c_{1}}\right)$.
@@
The assumption that $\delta\leq 1$ and $\epsilon\leq 1$ is used in the proof for Lemma 3 and has been made only for simplicity. These can be taken as any positive values, with changes required only in the constants $C$ and $D$.
@@
The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$. This can be bounded using high probability bounds for TD(0) (e.g., Korda and LA (2015)) can be used as well. Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al. 2018, Chen et al. 2021) can be combined with Markov’s inequality to bound this term.
@@ 3 Proof of the Main Result
where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-z^{*}\|$. To bound the second term, note that
@@
We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
@@
$x^{\prime}_{m+1}\leq(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
@@
$>1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$. Analogous notation is used for other random variables. For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}$ and $\bar{z}_{k,n-1}(\xi)=z_{k}$ for all $k$. For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$. Also, define $\bar{x}_{n,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq n}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$. Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
@@
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ $\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\leq P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
The second inequality here follows from the observation that
@@
which gives us the required inequality. The third inequality follows from union bound.
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$. Note that $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Using the definitions above,
@@
This completes the proof for Theorem 1.

### v2

@@ 1 Introduction
TD(0) is one of the most popular reinforcement learning algorithms for policy evaluation (Tsitsiklis and Van Roy 1997). Given a fixed policy, the algorithm is an iterative method to obtain the value function for each state under the long-term discounted reward framework. To mitigate the issues of large state spaces, the value function is often approximated using a linear combination of feature vectors. This algorithm is referred to as TD(0) with linear function approximation. In this paper, we work with online TD(0) with a single sample path of the underlying Markov chain. Our goal in this paper is to obtain a concentration bound of the form from some time on or more precisely, for all $n\geq n_{0}$ for a suitably chosen $n_{0}$ for this algorithm.
@@
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2023), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$). Korda and Prashanth (2015) give a high probability bound for the same online framework as us.
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2023), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
Inspired by Chen et al. (2020), one of us considered stochastic approximation involving contractive maps and martingale noise, and derived such concentration bounds ‘from some $n_{0}$ on’ for this class of algorithms (Borkar 2021). This covered in particular synchronous Q-learning for discounted cost and some related schemes. In Chandak et al. (2022), we extended this work to cover ‘Markov noise’ (Meerkov 1972) in the stochastic approximation scheme, allowing us to give bounds for the asynchronous case. As mentioned before, this work assumed almost-sure boundedness of iterates, which is not satisfied by the TD(0) algorithm. We remove the need for this assumption in our current work. Other articles aiming at bounds of these forms, such as the one we provide, are found in Chandak et al. (2023) for the LSPE algorithm, and Borkar (2002), Thoppe and Borkar (2019), Kamal (2010) for abstract stochastic approximation schemes. Another work which considers all-time bounds is Chen et al. (2023), but they only consider additive and multiplicative noise and not Markovian noise as considered in this paper.
@@ 2 TD(0) - Background and Main Result
The state space can often be large ($s\gg 1$), and to alleviate this ‘curse of dimensionality’, $V$ is often approximated using a linear combination of $d$ linearly independent basis functions (feature vectors) $\phi_{i}:S\mapsto\mathcal{R},1\leq i\leq d$, with $s>>d\geq 1$. Also, let $\varphi(i)=[\phi_{1}(i),\ldots,\phi_{d}(i)]^{T}$. Thus $V(i)\approx\sum_{j=1}^{d}x(j)\phi_{j}(i)$, i.e., $V(i)=x^{T}\varphi(i)$ and $V\approx\Phi x$ where $x=[x(1),...,x(d)]^{T}$ and $\Phi$ is an $s\times d$ matrix whose $i$th column is $\phi_{i}$. Here $x$ denotes the learnable weights for the linear function approximator. Since $\{\phi_{i}\}$ are linearly independent, $\Phi$ is full rank. Substituting this approximation into the dynamic programming equation above leads to
@@
    First we make an assumption on $\Phi$. For that, define $\Psi\coloneqq\Phi^{T}\sqrt{D}$ and let $\lambda_{M}$ be the largest singular value of $\Psi$, i.e., the square of the largest eigenvalue of $\Psi\Psi^{T}$ and equivalently, of $\Psi^{T}\Psi$. Assume that
@@
    Since the feature vectors can be scaled without affecting the algorithm (the weights $x(i)$ get scaled accordingly), this assumption does not restrict the algorithm. Alternatively, the stepsize can be appropriately scaled as well, i.e., $a(n)=b(n)c$, where $b(n)$ acts as the effective stepsize and $c\varphi(\cdot)$ act as the effective feature vectors which satisfy the above assumption. This assumption can be replaced with the following assumption on the basis vectors.
    $\|\varphi(i)\|\leq\frac{\sqrt{2(1-\gamma)}}{1+\gamma}\;\forall i,$
    i.e., the $\ell_{2}$ norm of each row of $\Phi$ is bounded by $\sqrt{2(1-\gamma)}/1+\gamma$. To see this, note that
    $\frac{\|\Psi x\|_{2}}{\|x\|_{2}}=\frac{\|\Phi x\|_{D}}{\|x\|_{2}}\leq\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}.$
    Now $\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}$ is the operator norm defined with $\ell_{2}$ norm for domain and $\ell_{\infty}$ for co-domain, which is equal to the maximum $\ell_{2}$ norm of a row. Hence
    $\lambda_{M}=\max_{x\neq\theta}\frac{\|\Psi x\|_{2}}{\|x\|_{2}}\leq\max_{x\neq\theta}\frac{\|\Phi x\|_{\infty}}{\|x\|_{2}}=\max_{i}\|\varphi(i)\|_{2}.$
@@
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor. Prior works such as Tsitsiklis and Van Roy (1997) show the contractive nature with respect to the weighted norm $\|\cdot\|_{D}$.
@@
$\alpha=\sqrt{1-\min_{x\neq\theta}\frac{\|\Phi x\|_{D}^{2}}{\|x\|^{2}}\bigg{(}2(1-\gamma)-\lambda_{M}^{2}(1+\gamma)^{2}\bigg{)}}.$
@@
The proof appears in Appendix B. The Banach contraction mapping theorem implies that $\sum_{i}\pi(i)F(\cdot,i)$ has a unique fixed point $x^{*}$, i.e., there exists a unique point $x^{*}$ such that $\sum_{i}\pi(i)F(x^{*},i)=x^{*}$. Then,
@@
    $1-2d\sum_{m=n_{0}+1}^{n}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
@@
    $1-2d\sum_{m\geq n_{0}+1}e^{-D\delta^{2}/\beta_{n_{0}}(m)}-P(\|x_{n_{0}}-x^{*}\|>\epsilon)$
@@
The assumption that $\delta\leq 1$ and $\epsilon\leq 1$ is used in the proof for Lemma 3 and has been made only for simplicity. These can be taken as any positive values, with changes required only in the constant $D$.
@@
The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$. This can be bounded using high probability bounds for TD(0) (e.g., Korda and Prashanth (2015)). Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al. 2018, Chen et al. 2021) can be combined with Markov’s inequality to bound this term.
Remark 3.
For the special case of $a(n)=\frac{d_{1}}{n}$, and by combining our result with the finite time bound in Prashanth et al. (2021), we can rewrite the result in Theorem 1 in the following form.
Corollary 1.
Let $a(n)=d_{1}/n$ with sufficiently large $d_{1}$, and $n_{0}$ large enough to satisfy assumptions of Theorem 1. Then with probability at least $1-\varepsilon$, we have for all $m\geq n_{0}$ that
$\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
The proof for this corollary has been presented at the end of Appendix B. Chen et al. (2023) gives an all-time bound which matches our result at $n_{0}$, but their bound goes to zero as $m\uparrow\infty$. In our bound, the term $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ remains constant as $m$ is increased. Here $\delta$ can be modified to $\delta(m)$ (similar to the treatment in Corollary 1 in Chandak et al. (2022)). But the term $a(n_{0})(c_{2}+c_{1}\epsilon)$ arises from our treatment of Markov noise using Poisson equation. Note that Chen et al. (2023) do not consider the case of Markov noise, but only consider the case of additive and multiplicative noise (i.i.d. samples of state-action-next state triplets). We leave incorporating their ideas into our approach to get a bound decaying with $m$ for Markov noise as a future work.
@@ 3 Proof of the Main Result
where $z_{n_{0}}=x_{n_{0}}$. Note that $\|x_{n}-x^{*}\|\leq\|x_{n}-z_{n}\|+\|z_{n}-x^{*}\|$. To bound the second term, note that
@@
Here we use the definition that $x_{n_{0}}=z_{n_{0}}$. We first simplify the third term above. For simplicity, we define $F(x,Y)=F_{1}(Y)+F_{2}(Y)x+x,$ where
@@
$x^{\prime}_{m+1}\leq$ $\;(\alpha+a(n_{0})c_{1})x^{\prime}_{m}+\left\|\sum_{k=n_{0}}^{m}\chi(m,k+1)a(k)\left(M_{k+1}x_{k}+\widetilde{W}_{k+1}x_{k}+\widetilde{U}_{k+1}\right)\right\|$
$+a(n_{0})(c_{2}+c_{1}\|x_{n_{0}}-x^{*}\|).$
@@
$\|x_{n_{0}}-x^{*}\|\leq\epsilon\implies\|z_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\|x_{n_{0}}-x^{*}\|.$
Also recall that $\sup_{n_{0}\leq m\leq n}\|x_{m}-z_{m}\|=x_{n}^{\prime}$. Hence
$\left\{\|x_{n_{0}}-x^{*}\|\leq\epsilon\right\}\bigcap\left\{x^{\prime}_{n}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right\}$
$\subseteq\left\{\|x_{m}-x^{*}\|\leq\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
This implies the following relation between the probabilities of the two sets.
@@
$\geq 1-P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\bigcup\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
@@
Here the notation $x_{m}^{\prime}(\xi)$ and $x_{n_{0}}(\xi)$ highlights the dependence of $x_{m}^{\prime}$ and $x_{n_{0}}$ on the realizations of $x_{0}$ and $\{Y_{k}\}$. Analogous notation is used for other random variables. For $\xi\notin\mathbb{B}_{n-1}$ let us define $\bar{x}_{k,n-1}(\xi)=x_{k}(\xi)$ and $\bar{z}_{k,n-1}(\xi)=z_{k}(\xi)$ for all $k$. For $\xi\in\mathbb{B}_{n-1}$, we define $\bar{x}_{k,n-1}(\xi)=x^{*}$, and $\bar{z}_{k,n-1}(\xi)=x^{*}$ for all $k$. Also, define $\bar{x}_{m,n-1}^{\prime}(\xi)=\sup_{n_{0}\leq k\leq m}\|\bar{x}_{k,n-1}(\xi)-\bar{z}_{k,n-1}(\xi)\|$. Now, adapting the proof of Proposition 34 from Tao and Vu (2015) for our case, we first note that $\xi\notin\mathbb{B}_{n-1}\implies\bar{x}_{n,n-1}^{\prime}(\xi)=x_{n}^{\prime}(\xi)$, which implies
@@
$P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$
$\leq P\left(x^{\prime}_{n}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\bigcup\;\bar{x}^{\prime}_{n,n-1}\neq x^{\prime}_{n}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\;\bigcup\;\|x_{n_{0}}-x^{*}\|>\epsilon\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P(\mathbb{B}_{n-1})$
$=P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)+P\left(x_{n-1}^{\prime}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\;\;\bigcup\;\;\|x_{n_{0}}-x^{*}\|>\epsilon\right).$
Inequality (a) here follows from the observation that
@@
which gives us the required inequality. Inequality (b) follows from union bound and inequality (c) follows from the observations that $\{\|x_{n_{0}}-x^{*}\|>\epsilon\}\subseteq\mathbb{B}_{n-1}$ and $\{\bar{x}_{n,n-1}^{\prime}\neq x_{n}^{\prime}\}\subseteq\mathbb{B}_{n-1}$.
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction. Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ for $n_{0}\leq m\leq n-1$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$. Then,
@@
This completes the proof for the first part of Theorem 1.
Let $A_{n}$ be the set
$\left\{\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall n_{0}\leq m\leq n\right\}.$
Then $\{A_{n}\}$ is a decreasing sequence of sets, i.e., $A_{n+1}\subseteq A_{n}$ for all $n\geq n_{0}$. Now let $A$ be the set
$\left\{\|x_{m}-x^{*}\|\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}},\;\forall m\geq n_{0}\right\}.$
Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1.